OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find 2·f(4).

OpenStudy (anonymous):

f(4) = 4^2 + 1 = 17

OpenStudy (anonymous):

2 . f(4) = ?

OpenStudy (anonymous):

2 * 17 = 34